Question content area top
Part 1
A company will need ​$75,000 in 7 years for a new addition. To meet this​ goal, the company deposits money in an account today that pays ​4% interest compounded annually. Find the amount to the nearest hundred dollars that should be invested to total ​$75,000 in 7 years.

Answer by math_tutor2020(3817)


Answer: $57,000


Work Shown:
A = P(1+r/n)^(n*t)
75,000 = P(1+0.04/1)^(1*7)
75,000 = P(1.31593177923584)
P = 75,000/(1.31593177923584)
P = 56,993.8359901548
P = 56,993.84
P = 57,000

The 993.84 bumps up to 1000 when rounding to the nearest hundred dollars, which leads to 56000+1000 = 57000
In other words, 56993.84 is closer to 57000 than it is to 56900.